# Context Matters: Peer-Aware Student Behavioral Engagement Measurement via VLM Action Parsing and LLM Sequence Classification

### v1

@@ 3.3 Student action recognition
We investigated few-shot learning for action recognition using vision-language models (VLMs). In particular, we use Video Fine-tuned CLIP (ViFi-CLIP) [23] because of its efficiency and simplicity. The ViFi-CLIP adapted CLIP to the video domain through fine-tuning both the CLIP image and text encoders. For a video segment $S_{i}\in\mathbb{R}^{TxHxWxC}$, the embeddings of its T frames are obtained using the CLIP image encoder and then average-pooled to get the video-level representation $v_{i}\in\mathbb{R}^{D}$ . The corresponding class label $C_{i}$ is embedded in a prompt (e.g., ‘a photo of a <<category>>’) and passed through the CLIP text encoder to produce a text embedding $t\in\mathbb{R}^{D}$. We fine-tuned the ViFi-CLIP model on the student actions dataset in a few-shot setting with 16 video clips for each action category by maximizing the cosine similarity between a batch of video-level embeddings $v_{i}$ and their corresponding text embeddings $t_{i}$ using Cross Entropy loss with entropy regularization,
@@ 3.4 Student engagement classification
To determine whether a student is engaged or disengaged, we leverage the LLM as a classifier in a zero-shot prompting. Following the prompt-based learning paradigm, the LLM is conditioned on a prompt $x_{prompt}$ to generate a predefined textual label $y\in Y$ (e.g., engaged, disengaged). The prompt $x_{prompt}$ is constructed by concatenating a task description $x_{desc}$ and input $x_{input}$ as $x_{desc}$$x_{input}$}. An example of the task description $x_{desc}$, which describes the engagement classification based on the student’s actions and classroom context, is shown in Fig. 6 An example of the input $x_{input}$, which includes the student action sequence with the duration of each action, along with the classroom context, represented in the majority of peers’ actions during the time interval, can be represented as:
@@ 4.1 Experimental Setup
Benchmarks. The lack of a public benchmark for classroom engagement analysis makes it necessary to create one. So, we created a behavioral engagement dataset with two subsets of annotated videos—one for student actions and the other for student engagement—in order to independently assess each step of the proposed framework. All videos were recorded from a wall-mounted 4K camera at 15 fps in classrooms of 6–11 students. For the student action dataset, we recorded 7 subjects performing 13 classroom-related actions in a controlled environment. This resulted in 208 trimmed clips for training (16 per action) and 47 trimmed video clips for testing across the 13 action categories. This trimmed video dataset is designed to evaluate the proposed action recognition stage (Stage 1). The student engagement dataset consists of videos from three half-hour lectures involving eleven students. Experts in psychology and education labeled each two-minute clip as either engaged or disengaged. Each two-minute video is also temporally partitioned, and every resulting segment is assigned a predefined action label. This untrimmed video dataset is used to evaluate the remaining stages: proposed temporal segmentation (Stage 2) and engagement classification (Stage 3).
Implementation details. We implemented the proposed framework in PyTorch on a workstation with two NVIDIA TITAN RTX GPUs (24 GB each). The VIFI-CLIP model (ViT-B/16 backbone) is fine-tuned in a few-shot setting with K=16 samples per class, using a batch size of 32 and a learning rate $2\times 10^{-6}$ for 50 epochs. We use the AdamW optimizer with a weight decay of 0.001 and apply cosine annealing for learning rate scheduling. Each video clip is represented by 32 uniformly sampled frames, which are resized to $224\times 224$. For the LLM, which classifies the student’s action sequence and classroom context as engaged or disengaged, we set the temperature to 0.1 to ensure a deterministic and non-creative output.
@@ 4.2 Evaluation of the proposed framework stages
Stage #1: Action Recognition. We evaluated state-of-the-art video-language models, including XCLIP, TC-CLIP, and ViFi-CLIP on our student classroom action test set (containing 48 instances) under zero-shot and few-shot protocols. For the zero-shot setting, the ViFi-CLIP achieved 52.1% top-1 accuracy, which is far from satisfactory. We attribute this poor performance to the model’s fine-tuning on general web videos (Kinetics-400). While this bridges the image-video gap for general tasks, it fails to capture the subtle, fine-grained differences in our domain, such as differences between reading and writing. For the few-shot setting, as shown in Tab. 1, we note that the performance improves for all methods as the number of shots increases. Across most values of K, ViFi-CLIP achieves the highest accuracy, with particularly noticeable gains in the low- and mid-shot regimes, demonstrating that it can be effectively adapted to our specialized domain with minimal data. For example, ViFi-CLIP achieves gains of 2.1%, 4.2%, and 2.1%, compared to XCLIP at K=4, K=8, and K=16, respectively.
@@
TC-CLIP [18] 43.8% 43.8% 77.1% 91.7%
XCLIP [22] 29.2% 72.9% 87.5% 95.8%
ViFi-CLIP [23] 39.6% 75.0% 91.7% 97.9%
Stage #2: Temporal action segmentation. Tab. 2 shows the comparison of Gemini 2.5 Pro, which is a multimodal large language model, and Gemini 2.5 flash with the proposed VLM action parsing on the student engagement dataset, which provides dense action labels for each 2-minute video. We curated a prompt describing the task that should be performed on the given 2-min video as shown in Fig. 8.
Figure 8: The prompt for temporal action segmentation.
It is noticeable from Tab. 2 that Gemini-2.5-pro achieves higher MoF, edit, and F1 scores compared to both Gemini-2.5-flash and the proposed VLM action parsing, while the proposed VLM action parsing outperforms Gemini-2.5-flash across all three metrics. Dense student action labeling is challenging due to fine-grained, visually similar actions (e.g., ‘looking at laptop’ vs. ‘looking down’) and cluttered environments. The task is further complicated by inter-student occlusion, where objects from one student’s space appear in another’s, creating ambiguous visual data. However, when we combined the visually similar actions e.g., writing on a notebook and reading, the performance is enhanced.
Table 2: Comparison of temporal action segmentation. Writing notes/reading means the two are merged into one class.
Method F1@10 F1@25 F1@50 Edit MoF
Gemini-2.5-pro 56.1 53.0 42.8 49.8 67.1
Gemini-2.5-flash 37.4 32.9 23.5 36.4 54.4
Our action parsing 44.6 39.3 27.1 42.0 58.9
Our action parsing (writing notes/Reading) 47.1 42.2 29.5 44.3 62.2
@@
To quantify the performance of the context-based engagement classifier, we first evaluate LLMs using the manual dense student action annotation, as shown in Tab. 3. We can notice that both Llama-3-8B and Gemma-2-9B achieve similar performance in terms of F1 score. The moderate performance of the LLMs on the disengaged class can be attributed to the inherent ambiguity of student engagement. In our dataset, some students were doodling on tablets rather than taking notes; experts labeled these cases as disengaged, but our action labels do not distinguish fine-grained behaviors such as writing versus drawing. Additionally, students who were daydreaming were sometimes annotated as listening because their posture suggested attention despite their lack of focus. Finally, student seated at the classroom periphery may appear to look forward without actually paying attention to the instructor. The context-aware engagement classifier achieves 92% F1 score when using ground truth action annotations. When employing automated action parsing, it attains 86% with Gemini-based parsing, Tab. 5, and 85% with the VLM-based parsing, Tab. 4. This performance gap is attributable to the difficulty of the action recognition task, as even the advanced Gemini-2.5-Pro model struggles to accurately parse the 2-minute-long student videos.
Table 3: Comparison of LLMs for context-based engagement classification based on manual annotation of student tubelets.
@@
Table 4: Comparison of LLMs for context-based engagement classification based on VLM action parsing of student tubelets.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.55 0.39 0.45
engaged 0.88 0.93 0.91
weighted avg 0.84 0.87 0.85
Gemma-2-9B disengaged 0.78 0.36 0.49
engaged 0.81 0.96 0.88
weighted avg 0.80 0.89 0.83
GPT3.5 disengaged 0.76 0.32 0.45
engaged 0.78 0.96 0.86
weighted avg 0.78 0.88 0.81
Table 5: Comparison of LLMs for context-based engagement classification based on Gemini-2.5-Pro-based temporal action segmentation of student tubelets.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.58 0.41 0.48
engaged 0.88 0.94 0.91
weighted avg 0.85 0.87 0.86
Gemma-2-9B disengaged 0.69 0.37 0.48
engaged 0.84 0.95 0.89
weighted avg 0.82 0.88 0.84
GPT3.5 disengaged 0.73 0.38 0.50
engaged 0.96 0.84 0.89
weighted avg 0.83 0.89 0.85
@@ 4.3 Ablation Studies
To ablate our model, we performed two experiments: a) we removed the classroom context to measure its contribution, and b) we compared our sequence-based method against a histogram-based counterpart. In this section, all engagement classification was performed by Llama-3-8B, using the ground-truth dense action labels for each 2-minute student video.
a) Engagement classification without classroom context: Tabs. 3 and 6 shows a comparison between the proposed peer-aware approach and its peer-free counterpart, respectively. It is noticeable that the proposed peer-aware approach outperforms its peer-free counterpart, achieving a 2% higher F1 score for the disengaged class. The prompt for engagement classification based on the student’s actions is shown in Fig. 10.
@@
Table 6: Comparison of context-free engagement classifier based on manual annotation of student tubelets using Llama-3-8B.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.75 0.51 0.61
engaged 0.90 0.96 0.93
weighted avg 0.88 0.91 0.89
Gemma-2-9B disengaged 0.73 0.60 0.66
engaged 0.93 0.96 0.95
weighted avg 0.91 0.92 0.91
GPT-3.5 disengaged 0.75 0.44 0.55
engaged 0.87 0.96 0.91
weighted avg 0.85 0.90 0.87
b) Input representation: The sequence-based and histogram-based input representations of a student’s time-stamped actions are compared in Tab. 7. We note that although the proposed sequence-based method’s overall F1 scores are comparable to that of its histogram-based counterpart, the sequence-based approach clearly outperforms the latter in terms of recall for the disengaged class (75% vs. 49%). This substantial difference supports our hypothesis that what really reflects student engagement is the temporal sequence of actions rather than an overall duration of an action represented by a histogram.
Table 7: Comparison of input representation for engagement classification based on manual annotation of student tubelets Using Llama-3-8B .
       Representation Recall Precision F1-score
Sequence disengaged 0.75 0.51 0.61
engaged 0.90 0.96 0.93
weighted avg 0.88 0.91 0.89
Histogram disengaged 0.49 0.77 0.60
engaged 0.98 0.93 0.96
weighted avg 0.92 0.91 0.91
5 Conclusion
We presented a novel context-aware framework to measure student behavioral engagement in a classroom setting. The proposed framework uses a Vision-Language Model (VLM) to parse student actions and a Large Language Model (LLM) to categorize engagement based on the resulting action sequence and the classroom context (i. e., peers’ activities). The experimental results demonstrated that context awareness is crucial to take into account because, depending on the context, the same action may indicate engagement or disengagement. On our created dataset, we evaluated the proposed framework and obtained F1 scores for engagement classification of 92%, 86%, and 85% with manual, Gemini-based, and modified VLM-based temporal action segmentation, respectively. Furthermore, the proposed framework overcomes the challenge of the scarcity of annotated data required to model the diversity of student behaviors.

### v2

@@ 3.3 Student action recognition
We investigated few-shot learning for action recognition using vision-language models (VLMs). In particular, we use Video Fine-tuned CLIP (ViFi-CLIP) [24] because of its efficiency and simplicity. For a video segment $S_{i}\in\mathbb{R}^{TxHxWxC}$, the embeddings of its T frames are obtained using the CLIP image encoder and then average-pooled to get the video-level representation $v_{i}\in\mathbb{R}^{D}$ . The corresponding class label $C_{i}$ is embedded in a prompt (e.g., ‘a photo of a <<category>>’) and passed through the CLIP text encoder to produce a text embedding $t\in\mathbb{R}^{D}$. We fine-tuned the ViFi-CLIP model on the student actions dataset in a few-shot setting with 16 video clips per action category by maximizing the cosine similarity between a batch of video-level embeddings $v_{i}$ and their corresponding text embeddings $t_{i}$ using cross-entropy loss. Additionally, we incorporate entropy regularization into the loss function to enhance generalization and mitigate overfitting by discouraging overly confident, incorrect predictions.
@@ 3.4 Student engagement classification
To determine whether a student is engaged or disengaged, we leverage the LLM as a classifier in zero-shot prompting. Following the prompt-based learning paradigm, the LLM is conditioned on a prompt $x_{prompt}$ to generate a predefined textual label $y\in Y$ (e.g., engaged, disengaged). The prompt $x_{prompt}$ is constructed by concatenating a task description $x_{desc}$ and input $x_{input}$ as $x_{desc}$$x_{input}$}. An example of the task description $x_{desc}$, which describes the engagement classification based on the student’s actions and classroom context, is shown in Sec. A.1.An example of the input $x_{input}$, which includes the student action sequence with the duration of each action, along with the classroom context, represented in the majority of peers’ actions during the time interval, can be represented as:
@@ 4.1 Experimental Setup
Benchmarks. The lack of a public benchmark for classroom engagement analysis makes it necessary to create one. So, we created a behavioral engagement dataset with two subsets of annotated videos—one for student actions and the other for student engagement—in order to independently assess each step of the proposed framework. All videos were recorded from a wall-mounted 4K camera at 15 fps in classrooms of 6–11 students. For the student action dataset, we recorded 7 subjects performing 13 classroom-related actions in a controlled environment. This resulted in 208 trimmed clips for training (16 per action) and 46 trimmed video clips for testing across the 13 action categories. This trimmed video dataset is designed to evaluate the proposed action recognition stage (Stage 1). The student engagement dataset comprises videos from three half-hour lectures involving eleven students. Each video is segmented into 2-minute clips per student, yielding a total of 455 video clips. Experts in psychology and education labeled each two-minute clip as either engaged or disengaged, resulting in 400 engaged and 55 disengaged samples. Each two-minute video segment is also temporally partitioned manually, and every resulting segment is assigned a predefined action label. The entire student engagement dataset is used as a test set to evaluate the remaining stages—proposed temporal action segmentation (Stage 2) and engagement classification (Stage 3)—since both stages are training-free: Stage 2 employs a temporal sliding window, while Stage 3 uses an LLM with zero-shot prompting.
Implementation details. We implemented the proposed framework in PyTorch on a workstation with two NVIDIA TITAN RTX GPUs (24 GB each). The VIFI-CLIP model (ViT-B/16 backbone) is fine-tuned in a few-shot setting with K=16 samples per class, using a batch size of 32 and a learning rate $2\times 10^{-6}$ for 50 epochs. We use the AdamW optimizer with a weight decay of 0.001 and apply cosine annealing for learning rate scheduling. Each video clip is represented by 16 uniformly sampled frames, which are resized to $224\times 224$. For an LLM-based engagement classifier, we set the temperature to 0.1 to ensure a deterministic and non-creative output.
@@ 4.2 Evaluation of the proposed framework stages
Stage #1: Action Recognition. We evaluated state-of-the-art video-language models, including XCLIP [23], TC-CLIP [19], and ViFi-CLIP [24] on our student classroom action test set (46 instances) under zero-shot and few-shot protocols. For the zero-shot setting, the ViFi-CLIP achieved 52.1% top-1 accuracy, which is far from satisfactory. We attribute this poor performance to the model’s fine-tuning on general web videos (Kinetics-400). While this bridges the image-video gap for general tasks, it fails to capture the subtle, fine-grained differences in our domain, such as differences between reading and writing. For the few-shot setting, as shown in Tab. 1, we note that the performance improves for all methods as the number of shots increases. Across most values of K, ViFi-CLIP achieves the highest accuracy, with particularly noticeable gains in the low- and mid-shot regimes, demonstrating that it can be effectively adapted to our specialized domain with minimal data. For example, ViFi-CLIP achieves gains of 4.5%,2.2%, and 4.6%, compared to XCLIP at k=2, K=4, and K=8, respectively.
@@
TC-CLIP [19] 52.2% 52.2% 80.4% 93.5%
XCLIP [23] 43.2% 79.6% 88.6% 97.9%
ViFi-CLIP [24] 47.7% 81.8.0% 93.2% 97.9%
Stage #2: Temporal action segmentation (TAS). Tab. 2 shows the comparison of Gemini 2.5 Pro and Gemini 2.5 Flash, which are multimodal large language models (MLLMs), with the proposed VLM action parsing on the student engagement dataset, which provides dense action labels for each 2-minute video. For MLLMs, we curated a prompt describing the task to be performed on the given 2-min video as shown in Sec. A.2.
Dense student action labeling is challenging due to fine-grained, visually similar actions (e.g., ‘looking at laptop’ vs. ‘looking down’) and cluttered environments. The task is further complicated by inter-student occlusion, where objects from one student’s space appear in another’s, creating ambiguous visual data. It is noticeable from Tab. 2 that Gemini-2.5-pro achieves higher MoF, edit score, and F1 score compared to both Gemini-2.5-flash and the proposed VLM action parsing, while the proposed VLM action parsing outperforms Gemini-2.5-flash across all three metrics. Moreover, the proposed VLM action parsing is more consistent in identifying the on- and off-task actions than both Gemini MLLMs. This leads to accurate engagement classification, as shown next.
Table 2: Comparison of temporal action segmentation using VLM-based, Gemini-2.5-Flash, and Gemini-2.5-Pro-based.
TAS Accuracy Edit score F1@10 F1@25 F1@50
Gemini-2.5-flash 57.2 37.4 39.0 34.5 25.3
Gemini-2.5-pro 69.8 51.8 58.2 55.2 45.2
VLM-based 67.0 45.7 48.3 43.9 31.4
@@
To quantify the performance of the context-aware engagement classifier, we first evaluate LLMs using the manual dense student action annotation, as shown in Tab. 3. We can notice that both Llama-3-8B [13] and Gemma-2-9B [32] achieve similar performance in terms of F1 score. The moderate performance of the LLMs on the disengaged class can be attributed to the inherent ambiguity of student engagement. In our dataset, some students were doodling on tablets rather than taking notes; experts labeled these cases as disengaged, but our action labels do not distinguish fine-grained behaviors such as writing versus drawing. Additionally, students who were daydreaming were sometimes annotated as listening because their posture suggested attention despite their lack of focus. Finally, student seated at the classroom periphery may appear to look forward without actually paying attention to the instructor. Since the context-aware engagement classifier (Gemma) achieves the highest F1-score of 92% using manually annotated actions, we subsequently employ Gemma with automated action parsing. As shown in Tab. 4, it achieves an F1-score of 86% with Gemini-based parsing and 88% with VLM-based parsing. This performance gap is attributable to the difficulty of the student action recognition task, as even the advanced Gemini-2.5-Pro model struggles to accurately parse the 2-minute-long student videos.
Table 3: Comparison of LLMs for context-aware engagement classification based on manual action annotation of student tubelets.
@@
Table 4: Comparison of context-aware engagement classifier based on VLM action parsing and Gemini-2.5-pro-based temporal segmentation of student tubelets using Gemma-2-9B.
       Action parsing of student tubelet Recall Precision F1-score
Gemini-2.5-Pro-based disengaged 0.58 0.43 0.49
engaged 0.89 0.94 0.92
weighted avg 0.85 0.88 0.86
Our VLM-based disengaged 0.53 0.52 0.52
engaged 0.93 0.93 0.93
weighted avg 0.88 0.88 0.88
4.3 Comparison with Existing Works
Direct comparison with prior work is difficult due to differences in datasets, action classes, and annotation protocols. Therefore, we implemented the approach in [2] by leveraging the predicted actions from the fine-tuned VLM model and applying its reported best-performing classifier. For each student tublet, we obtain the sequence of predicted actions and construct the histogram of actions that represents the frequency of actions within a 2-minute interval. This histogram is then used as an input for the Random Forest classifier for engagement measurement. Since the method in [2] requires training, we evaluate it using Leave-One-Out Cross-Validation (LOOCV) on the entire dataset, which contains 3 30-minute lectures. As shown in Tab. 5, the proposed framework surpasses the compared method by 3% in F1-score. It also achieves a higher F1-score for the disengaged class (52% vs. 47%). This improvement highlights the effectiveness of our approach, which uses a student’s action sequence augmented with the actions of their peers.
Table 5: Comparison with existing work based on the proposed VLM-based action parsing of student tubelets using Gamma-2.9B.
       Method Recall Precision F1-score
[2] disengaged 0.70 0.36 0.47
engaged 0.84 0.97 0.90
weighted avg 0.83 0.91 0.85
Ours disengaged 0.53 0.52 0.52
engaged 0.93 0.93 0.93
weighted avg 0.88 0.88 0.88
4.4 Cross-Course Evaluation
Our student engagement dataset, consisting of 3 30-minute lectures, was collected from a single course. This raises a question: Can the proposed framework generalize to different classes with new students, instructors, and course content? To address this, we collected an additional 30-minute lecture from a separate class involving eight different students. This yielded 113 two-minute segments, consisting of 73 engaged and 40 disengaged samples. Tab. 6 presents a comparison of the context-aware engagement classifier based on ground truth manual annotation, Gemini-2.5-Pro-based temporal action segmentation (TAS), and our VLM-based TAS on a 30-minute lecture from a different class. The results show that the proposed framework with VLM-based TAS surpasses the Gemini-2.5-Pro-based variant by a substantial margin of 9% in weighted F1-score. This improvement highlights the generalization capability of the proposed VLM–LLM framework, demonstrating its ability to accurately parse and classify student action sequences despite variations in student cohorts and student–teacher dynamics.
Table 6: Evaluation of a context-aware engagement classifier on a 30-minute lecture from different class using Gemma-2-9B.
       Student tubelet annotation Recall Precision F1-score
Manual annotation disengaged 0.85 0.81 0.83
engaged 0.89 0.92 0.90
weighted avg 0.88 0.88 0.88
Gemini-2.5-Pro-based disengaged 0.90 0.56 0.69
engaged 0.62 0.92 0.74
weighted avg 0.72 0.79 0.72
Our VLM-based disengaged 0.65 0.84 0.73
engaged 0.93 0.83 0.88
weighted avg 0.83 0.83 0.83
@@ 4.5 Ablation Studies
To ablate our model, we conducted two experiments: a) removing the classroom context to assess its contribution, and b) comparing our sequence-based method with a histogram-based counterpart. In this section, all engagement classification experiments were performed using Gemma-2.9B, based on the action sequences obtained from our VLM-based action parsing for each 2-minute student video.
a) Engagement classification without classroom context: Tab. 7 shows a comparison between the proposed peer-aware approach and its peer-free counterpart, respectively. It is noticeable that the proposed peer-aware approach outperforms its peer-free counterpart, achieving a 1% higher F1-score for the disengaged class. Moreover, a context-free approach performs badly ($50\%$precision) in identifying disengaged students, which is critical. The prompt for context-free engagement classification is shown in Sec. A.1.
Table 7: Comparison of engagement classifier with/out classroom context based on the introduced VLM-based action parsing of student tubelets using Gemma-2-9B.
       w/o Recall Precision F1-score
Context-free disengaged 0.62 0.44 0.51
engaged 0.89 0.94 0.92
weighted avg 0.86 0.88 0.87
context-aware disengaged 0.53 0.52 0.52
engaged 0.93 0.93 0.93
weighted avg 0.88 0.88 0.88
b) Input representation: The sequence-based and histogram-based input representations of a student’s time-stamped actions are compared in Tab. 8. We note that the proposed sequence-based method achieved an 88% F1-score compared to 86% for its histogram-based counterpart. Specifically, the sequence-based approach clearly outperforms the latter in terms of F1-score for the disengaged class (52% vs. 39%). This substantial difference shows the student engagement is better captured by the temporal action sequence rather than by aggregated action durations represented as a histogram. Furthermore, temporal action sequences preserve the order and timing of events, enabling relational analysis of a student’s behavior in the context of peers’ concurrent actions. In contrast, action histograms lack temporal structure and therefore fail to capture such contextual dynamics.
Table 8: Comparison of input representation for engagement classification based on the introduced VLM action parsing of student tubelets Using Gemma-2-9B.
       Representation Recall Precision F1-score
Histogram disengaged 0.35 0.44 0.39
engaged 0.94 0.91 0.93
weighted avg 0.87 0.86 0.86
Sequence disengaged 0.53 0.52 0.52
engaged 0.93 0.93 0.93
weighted avg 0.88 0.88 0.88
5 Conclusion
We presented a novel context-aware framework to measure student behavioral engagement in a classroom setting. The proposed framework employs a Vision-Language Model (VLM)-based action parser to get a sequence of student actions during an arbitrary time interval and then feeds this student sequence, along with the classroom context (i.e., peers’ activities), to a Large Language Model (LLM) for engagement measurement. The experimental results demonstrated that context awareness is crucial to take into account to detect disengaged students because, depending on the context, the same action may indicate engagement or disengagement. On our constructed dataset, we evaluated the proposed framework and obtained engagement classification F1 scores of 92%, 86%, and 88% using manual, Gemini-based, and modified VLM-based temporal action segmentation, respectively. Furthermore, the proposed framework overcomes the challenge of the scarcity of annotated data required to model the diversity of student behaviors.
6 Ethical Statement
This study was approved by IRB # 19.0513 and informed consent was obtained from all participants.
Appendix A Appendix
A.1 LLM Prompts
In this section, we present the LLM prompts used for:
  • Context-free engagement classification (see Fig. 6).
  • Context-aware engagement classification (see Fig. 8).
The context-free engagement classification uses only a student’s action sequence as input and outputs the student’s engagement level along with the reasoning behind that assessment. In contrast, context-aware engagement classification takes both a student’s action sequence and the majority action sequence of their peers as input and outputs the student’s engagement level together with the reasoning for that assessment.
@@ A.1 LLM Prompts
Figure 8: An example of the task description $x_{desc}$
A.2 MLLM Prompt
To assess the capability of advanced multimodal large language models, e.g., Gemini-2.5 Pro, in parsing a student’s actions within a 2-minute video, we designed a prompt that instructs the MLLM to perform temporal action segmentation as shown in Fig. 10.
Figure 10: The prompt for temporal action segmentation.